Hi, can someone please help me solve this equation: it says ,
Solve the system by graphing:In order to receive credit for this problem you must show all your work.
The equation is
(y=2x+4)
(y=-x+7)
thank you in advance!

Found 2 solutions by drpipespropmzy, precise:
Answer by drpipespropmzy(2)

The equation is
(y=2x+4)
(y=-x+7)
They are lines and must be represented graphically.
For line one y=2x+4 and line two y=-x+7
For line 1 , intercept on x and y axis;
When y=0 it implies 0=2x+4
x=-2, hence (-2,0)
y=2(0)+4, (0,4)
For line 2, intercept on x and y axis;
When y=0 it implies 0=-x+7
x=7, hence (7,0)
When x=0 it implies y=-(0)+7, (0,7)
So the points on line 1 and 2 are
(-2,0) ,(0,4) and (7,0) , (0,7).
Now use these points to draw the graph whiles you've shown all the workings involved.

Answer by precise(6)

Okay, well let's take that first equation. Y=2x+4: the 2 is your slope. (Up two over 1.) And the 4 is where it crosses the y axis. So from the origin, go up four and put a point. From there go up two over 1 to your right for your next point, and so on. This is your slope. Don't forget to go down two and over 1 to your left from the point on four. Connect your dots. It should look like a straight slanted upward line.
Onto your next equation. Y=-x +7: the -1x is your slope. (Down 1 over 1 to your right.) The 7 is where it crosses the y axis. So from the origin, go up seven and put a point. From there, go down one over 1 to your right for your next point and so on.don't forget to go up one and over 1 to your left from the point on seven. This is your slope. Connect your dots . The line should be slanted downward.
Check your work. All together it should look like an x.
Sorry in advance if this was confusing. It really isn't , I promise. ;)